# Generalized Pinching-Antenna Systems: A Tutorial on Principles, Design Strategies, and Future Directions

### v1

@@ III-A1 Single Pinching Antenna on the Waveguide
In addition to the rate maximization problem, the transmit power minimization problem has also been investigated in [41] for NOMA transmission with two downlink users. Specifically, the authors of [41] addressed the challenge of jointly optimizing the pinching-antenna position and user power allocation to minimize the total transmit power while ensuring that individual user data rate requirements are satisfied. A closed-form analytical solution is derived, showing that the optimal pinching-antenna position typically favors the user closer to the waveguide. This strategic placement not only enhances the received signal quality for the weaker user but also facilitates efficient SIC. These results provide valuable theoretical insights and offer practical guidelines for deploying power-efficient NOMA in pinching-antenna systems.
@@
To compare different transmission schemes, Fig. 6 evaluates the average sum rate performance of NOMA and TDMA schemes in a two-user pinching-antenna system. In the simulation, we consider a square communication area with side length of $20$ m, i.e., $D_{x}=D_{y}=20$ m, and the pinching antenna is dynamically positioned to optimize the performance according to each scheme’s design. The SNR threshold for the primary user is set to $\gamma_{p}=1$. The results show that NOMA consistently outperforms TDMA across all considered transmit power levels. As the transmit power increases, the gain offered by NOMA becomes increasingly significant. This performance advantage is due to the fact that NOMA allows both users to be served simultaneously via power-domain multiplexing, thereby utilizing spectral resources more efficiently. The observed results clearly demonstrate the benefit of adopting NOMA-assisted pinching-antenna systems, especially in scenarios requiring high data rates and efficient resource utilization.
Figure 6: Average data rate of TDMA versus NOMA for a two-user pinching-antenna system.
@@
Research Insight: Multi-User Pinching-Antenna System Design Low-complexity TDMA Scheme: Repositioning the pinching antenna for each user in its allocated time slot achieves near-optimal performance based on analytical solutions, while a fixed antenna position offers reduced system complexity at the cost of a performance loss. Spectrally-Efficient NOMA Scheme: NOMA efficiently improves the system spectrum efficiency, where joint optimization of the antenna position and the user power allocation coefficients is crucial for balancing data rates and guaranteeing QoS constraints. Advanced Optimization Methods: For scenarios with more than two users or more complex system requirements, computationally intensive iterative algorithms are often needed if NOMA scheme is employed. Developing more efficient algorithms for these challenging optimization problems remains as an important and promising direction for future research.
@@ III-A2 Multiple Pinching Antennas on the Waveguide
$\underset{\tilde{x}_{1},...,\tilde{x}_{n}}{{\rm maximize}}$~ $\bigg|\sum\limits_{n=1}^{N}\frac{e^{-j\phi_{n}}}{(\tilde{x}_{n}-\bar{x})^{2}+C_{n}}\bigg|$ (17a)
$\phi_{n}=\frac{2\pi}{\lambda}[(\tilde{x}_{n}-\bar{x})^{2}+C]+\frac{2\pi}{\lambda_{g}}\tilde{x}_{n},\forall n\in{\mathcal{N}},$ (17c)
@@
To maximize the objective function, one needs to minimize the impact of the large scale path loss, and meanwhile, needs to ensure that the received signals from different pinching antennas are constructively combined at the user. To achieve this goal, one can solve the following relaxed problem:
@@
$h_{m}=\sum_{n=1}^{N}\frac{e^{-j\big(\frac{2\pi}{\lambda}[(\tilde{x}_{n}-\bar{x})^{2}+C_{m}]+\frac{2\pi}{\lambda_{g}}\tilde{x}_{n}\big)}}{\big[(\tilde{x}_{n}-\bar{x})^{2}+C_{m}\big]^{-\frac{1}{2}}},\forall m\in{\mathcal{M}},$ (19)
@@
Similar to the single-antenna setting, both TDMA and NOMA transmission can be employed. In the following, we discuss the design principles, resource allocation strategies, and associated challenges for each scheme.
@@
$R_{m}=t_{m}\log\left(1+\frac{\eta P_{\max}|h_{m}|^{2}}{N\sigma_{m}^{2}t_{m}}\right).$ (20)
@@
Besides alternating optimization with SCA, a recent study [57] has investigated the application of graph neural networks (GNNs) to address these challenges. In this framework, the objective is to jointly optimize the antenna placement and power allocation to maximize the system’s energy efficiency, subject to transmit power constraints and minimum inter-antenna spacing. To overcome the high computational complexity faced by conventional optimization methods, the GNN-based method treats the wireless system as a bipartite graph, with user and antenna nodes connected by edges that encode their spatial relationships. A bipartite GNN is then employed to learn effective mappings from the system state—comprising user and antenna positions as well as current channel conditions—to resource allocation decisions, specifically the antenna positions and the power allocated to each antenna. The learning process is supervised by training data generated from solutions obtained through conventional optimization techniques, enabling the GNN to approximate complex decision boundaries without requiring real-time, iterative optimization. Once trained, the GNN-based method can produce high-quality resource allocation policies with significantly lower computational overhead, making it attractive for practical systems with stringent latency requirements or large-scale user deployments.
@@
Problem (22) is challenging to solve due to the intricate coupling between the power allocation coefficients and the effective channel gains, which themselves are nonlinear functions of the pinching-antenna positions. To address this, a practical solution is to adopt a BCD approach, which decomposes the original problem into more tractable subproblems that can be solved iteratively [48]. Specifically, the optimization proceeds as follows:
  • Power allocation optimization: For given pinching-antenna positions, the optimal power allocation coefficients $(\alpha_{p},\alpha_{s})$ can be derived in closed-form based on the analysis in [48]. Notably, if the optimal pinching-antenna positions are available, e.g., by a linear search, this closed-form solution provides the globally optimal power allocation. This result offers a valuable theoretical benchmark for evaluating practical suboptimal algorithms with low computational complexity. However, exhaustive search quickly becomes infeasible as the number of antennas increases, motivating the need for efficient, scalable algorithms.
@@
Based on the BCD-SCA algorithm described above, we examine how the number of pinching antennas impacts the achievable data rate in the NOMA transmission scheme. The results are shown in in Fig. 7, where the parameters in Table IV is used and the two users are randomly deployed in a square region with side length $D=5$ m. As seen, the data rate of the secondary user increases monotonically as the number of pinching antennas grows. This improvement stems from the increased channel diversity and enhanced effective channel gain brought by each additional antenna, which collectively lead to stronger received signals and more effective interference management for the secondary user. As a result, deploying more pinching antennas along the waveguide directly contributes to improved spectral efficiency and higher achievable data rates. These findings underscore the practical importance of properly selecting and optimizing the number of pinching antennas in system design, as it enables significant performance gains while balancing hardware and implementation costs.
Figure 7: Achievable data rate of the secondary user versus different numbers of NOMA-assisted pinching-antenna systems with $\gamma_{p}=0.1$.
@@
Research Insights: Pinching-Antenna Systems with Multiple Pinching Antennas on One Waveguide Spatial Diversity Gains: Deploying multiple pinching antennas enables significant spatial diversity and potential array gains, which can be exploited to improve system performance. Constructive Signal Reception: Unlike conventional MIMO systems, beamforming is not feasible as only one signal can be injected into the waveguide at a time. Instead, optimizing the pinching-antenna positions to ensure that the received signals from all antennas are constructively combined at the user is crucial for performance improvement. System Design Challenge: In the multi-antenna scenario, the system design problem becomes considerably more complex, as it requires the joint optimization of pinching-antenna positions and resource allocation. Developing scalable, low-complexity algorithms for these highly coupled, non-convex problems, especially for large-scale or real-time systems, remains an important and open research direction.
@@ III-B Multiple-Waveguide Systems
To further enhance system flexibility and spatial diversity, multiple waveguides can be deployed in pinching-antenna systems. In this section, we systematically discuss design approaches for such multi-waveguide configurations by focusing on two representative scenarios. The first scenario assumes that each waveguide is equipped with a single pinching antenna, while the second one explores the use of multiple pinching antennas per waveguide. In the remainder of this subsection, we present the signal models, formulate the corresponding optimization problems, and highlight representative solutions for each case.
@@ III-B1 Single Pinching Antenna Per Waveguide
where $e_{m}$ denotes the mean square error of estimating the information of user $m$, while ${\bf w}=[w_{1},\ldots,w_{M}]^{\top}$ are the minimum mean square error (MMSE) receivers and weights, respectively. Compared to problem (24), the WMMSE reformulation offers a more tractable structure: the objective function is convex with respect to the beamforming vectors ${\bf v}_{m}$ and the MMSE weights, though not necessarily in the antenna positions $\tilde{x}_{n}$. In contrast, the original problem contains fractional forms involving the channel vectors and beamformers, leading to a highly coupled and challenging optimization problem.
Because the optimization variables are decoupled in the constraints, the WMMSE problem can be iteratively solved by the BCD approach. Nevertheless, as highlighted in [36], optimizing the pinching-antenna positions remains non-convex, and a linear search is typically needed, which can result in substantial computational overhead for systems with long waveguides or high-frequency operation. To mitigate this issue, a low-complexity two-stage algorithm was proposed in [36]. In the first stage, an MRT-based method is used to efficiently determine the pinching antenna positions. With these locations fixed, the second stage applies the standard WMMSE algorithm to optimize the beamforming matrix. This two-stage strategy circumvents the need to perform a linear search over the antenna positions in each iteration of the WMMSE algorithm, thereby greatly reducing the overall computational complexity while maintaining near-optimal performance.
@@
This setting is more challenging than the standard sum-rate maximization problem, due to the introduction of additional QoS requirements and more complex coupling between variables. Traditional BCD-type methods can become computationally prohibitive as network size and constraint set grow. To address these limitations, a gradient-based meta-learning joint optimization algorithm was proposed in [56]. The core idea is to “learn to optimize”: two neural networks are trained, one to update the beamforming vectors ${\bf v}_{m}$} and another one to update the pinching-antenna positions $\tilde{{\bf x}}$, by mimicking a gradient-based optimization process across a diverse set of channel realizations and QoS scenarios. During training, the networks are optimized to minimize the average negative weighted sum-rate, with the power and minimum rate constraints incorporated via penalty terms in the loss function. Once trained, the meta-learner can rapidly produce near-optimal beamforming and antenna placement solutions for new, unseen system configurations, yielding significant improvements in computational efficiency and scalability. Numerical results demonstrate that the meta-learning approach achieves superior weighted sum-rate performance and rapid convergence compared to classical methods, particularly in scenarios with stringent QoS requirements. These findings indicate that data-driven and meta-learning-based algorithms hold strong promise for tackling complex, large-scale joint optimization problems for pinching-antenna systems.
@@ III-B2 Multiple Pinching Antennas Per Waveguide
$h_{p,m}=\sum_{n=1}^{N}\frac{e^{-j\big(\frac{2\pi}{\lambda}[(\tilde{x}_{p,n}-\bar{x})^{2}+C_{m,p}]+\frac{2\pi}{\lambda_{g}}\tilde{x}_{p,n}\big)}}{\big[(\tilde{x}_{p,n}-\bar{x})^{2}+C_{m,p}\big]^{-\frac{1}{2}}},$ (27)
@@
Based on the channel model in (27), several recent studies have investigated pinching-antenna-enabled MIMO communications [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. Specifically, the authors of [67] investigated energy-efficient pinching-antenna system designs by jointly optimizing the pinching-antenna positions and downlink beamformers. To solve the challenging optimization problem, an iterative algorithm was proposed based the BCD and SCA methods. The study demonstrates substantial transmit power savings compared to conventional massive MIMO systems. A staged GNN architecture has been proposed in [69], which efficiently learns both antenna placement and beamforming, enabling scalable and accurate resource allocation for pinching-antenna systems with strong generalization across diverse network sizes and user distributions. Meanwhile, the authors of [70] introduced a Karush-Kuhn-Tucker (KKT)-guided dual learning framework based on a transformer architecture, enabling rapid and high-quality joint optimization of transmit and pinching beamforming in multi-user pinching-antenna systems. The authors of [79] explored the use of pinching-antenna systems for uplink and downlink multiuser MIMO communication. The digital precoder and pinching-antenna positions are jointly optimized to maximize the weighted sum-rate. Two low-complexity iterative algorithms are proposed based on fractional programming (FP) and Gauss-Seidel methods. The results show that pinching-antenna systems offer significant throughput gains over conventional massive MIMO and hybrid analog-digital systems, highlighting their promise for next-generation wireless networks. A recent study [80] investigated a practical implementation of pinching-antenna systems in multi-waveguide scenarios, where antennas are installed at pre-configured discrete locations rather than being continuously adjustable. Focusing on downlink NOMA transmission, the authors formulated a joint optimization problem involving waveguide assignment, antenna activation, SIC decoding order, and power allocation. By modeling assignment and activation as coalition-formation games and applying monotonic optimization and SCA, both globally optimal and near-optimal low-complexity solutions are obtained.
@@
Research Insights: Pinching-Antenna Systems with Multiple Waveguides Enhanced Spatial Multiplexing: Leveraging multiple waveguides provides greater degrees of freedom for pinching antenna deployment, substantially boosting spatial diversity, user capacity, and spectral efficiency compared to single-waveguide or fixed-antenna systems. System Design Challenges and Solutions: Achieving optimal performance requires joint optimization of pinching-antenna positions and beamforming, resulting in highly coupled and non-convex problems. Iterative algorithms, such as BCD and WMMSE/FP-based methods, are effective in tackling these challenges. Nevertheless, developing more practical and scalable algorithms for real-time and large-scale deployment remains an important research direction. Pinching Antenna Activation: In addition to pinching-antenna optimization, efficient antenna activation, i.e., activating pinching antennas at pre-configured positions is a practical and important research direction, which enables flexible system deployment while reducing implementation complexity.
III-C Uplink Pinching-Antenna Systems
Recent literature has begun to address these challenges from several perspectives [53, 50, 79, 83, 84]. The first detailed study of the multi-user uplink in pinching-antenna systems was [53], where TDMA was used to avoid inter-channel interference. The authors formulated and solved an joint optimization problem to determine the radio resources allocation and pinching-antenna positions optimization to maximize the minimum data rate among users, thereby ensuring fairness and meeting diverse QoS requirements. To further improve spectral efficiency, NOMA-based uplink pinching-antenna systems were studied in [50]. Considering a MIMO setting, the authors of [79] investigated the joint design of both uplink and downlink transmission for pinching-antenna systems. The sum-rate maximization problem was formulated and solved using an iterative algorithm. Simulation results clearly demonstrated the performance gain of pinching-antenna systems in uplink MIMO scenarios as compared to conventional fixed-antenna systems.
To address this issue, a recent study [84] proposed a segmented waveguide-enabled pinching-antenna architecture. In this design, the traditional long dielectric waveguide is divided into multiple shorter segments, each equipped with its own feed point and associated pinching antennas. By ensuring that only one pinching antenna is active per segment, this approach effectively eliminates inter-antenna radiation leakage and enables a tractable, physically consistent uplink signal model that supports multi-antenna operation. The work further introduced three practical operating protocols, including the segment selection, the segment aggregation, and the segment multiplexing, each offering a distinct trade-off in terms of performance and implementation complexity.
Table V provides a structured summary of representative system design problems and corresponding design strategies for pinching-antenna systems. It highlights the diversity of architectures, optimization variables, and transmission schemes addressed in the literature. This comparative overview underscores the adaptability of the pinching-antenna framework, the progression from closed-form designs to advanced iterative and learning-based algorithms, and the importance of system complexity and application context in shaping the choice of solution strategy. These foundations set the stage for addressing more challenging and practical issues in future research on generalized pinching-antenna systems.
TABLE V: Summary of representative system design problems and solutions for pinching-antenna systems.

### v2

@@ III-A1 Single Pinching Antenna on the Waveguide
In addition to the rate maximization problem, the transmit power minimization problem has also been investigated in [46] for NOMA transmission with two downlink users. Specifically, the authors of [46] addressed the challenge of jointly optimizing the pinching-antenna position and user power allocation to minimize the total transmit power while ensuring that individual user data rate requirements are satisfied. A closed-form analytical solution is derived, showing that the optimal pinching-antenna position typically favors the user closer to the waveguide. This strategic placement not only enhances the received signal quality for the weaker user but also facilitates efficient SIC. These results provide valuable theoretical insights and offer practical guidelines for deploying NOMA in pinching-antenna systems.
@@
To compare different transmission schemes, Fig. 7 evaluates the average sum rate performance of NOMA and TDMA for a two-user pinching-antenna system. We consider a square communication area with side length equals to $20$ m, i.e., $D_{x}=D_{y}=20$ m. The pinching antenna is dynamically positioned according to the design of each scheme. For the TDMA configuration, the transmission frame is divided into two equal-length time slots. Each user is served exclusively in one slot. Within its slot, the pinching-antenna position and transmit power are optimized for that user. For NOMA, both users are served simultaneously in the same time-frequency resource via power-domain multiplexing, with the pinching-antenna position and power allocation jointly optimized. The SNR threshold for the primary user $\gamma_{p}$ is set to $1$. The results show that NOMA consistently outperforms TDMA across the considered transmit power range, and the performance gap increases as the transmit power grows. This gain stems from the more efficient use of spectral resources enabled by NOMA, which allows both users to be served simultaneously.
We note that the considered design principles can be extended to scenarios with more than two users by combining pinching-antenna optimization with TDMA/NOMA-based resource allocation.
Figure 7: Average sum rate of TDMA versus NOMA for a two-user pinching-antenna system with different transmit power.
@@
Low-complexity TDMA Scheme: Repositioning the pinching antenna for each user in its allocated time slot achieves near-optimal performance based on analytical solutions, while a fixed antenna position offers reduced system complexity at the cost of a performance loss. Spectrally-Efficient NOMA Scheme: NOMA efficiently improves the system spectrum efficiency, where joint optimization of the antenna position and the user power allocation coefficients is crucial for balancing data rates and guaranteeing QoS constraints. Advanced Optimization Methods: For scenarios with more than two users or more complex system requirements, computationally intensive iterative algorithms are often needed if NOMA scheme is employed. Developing more efficient algorithms for these challenging optimization problems remains as an important and promising direction for future research.
@@ III-A2 Multiple Pinching Antennas on the Waveguide
$\underset{\tilde{x}_{1},...,\tilde{x}_{n}}{{\rm maximize}}$~ $\bigg|\sum\limits_{n=1}^{N}\frac{e^{-j\phi_{n}}}{(\tilde{x}_{n}-\bar{x})^{2}+C}\bigg|$ (17a)
$\phi_{n}=\frac{2\pi}{\lambda}[(\tilde{x}_{n}-\bar{x})^{2}+C]+\frac{2\pi}{\lambda_{g}}\tilde{x}_{n},\forall n\in{\mathcal{N}},$ (17d)
@@
To maximize the objective function, one needs to minimize the impact of the large scale path loss, and meanwhile, needs to ensure that the received signals from different pinching antennas are constructively combined at the user. To achieve this goal, one can solve the following approximated problem:
@@
$h_{m}=\sum_{n=1}^{N}\frac{\eta^{\tfrac{1}{2}}e^{-j\big(\frac{2\pi}{\lambda}[(\tilde{x}_{n}-\bar{x})^{2}+C_{m}]+\frac{2\pi}{\lambda_{g}}\tilde{x}_{n}\big)}}{\big[(\tilde{x}_{n}-\bar{x})^{2}+C_{m}\big]^{-\frac{1}{2}}},\forall m\in{\mathcal{M}},$ (19)
@@
Similar to the single-antenna setting, both TDMA and NOMA transmission can be employed. In the following, we discuss the design principles, resource allocation strategies, and the associated challenges for each scheme.
@@
$R_{m}=t_{m}\log\left(1+\frac{P_{\max}|h_{m}|^{2}}{N\sigma_{m}^{2}t_{m}}\right).$ (20)
@@
Besides alternating optimization with SCA, a recent study [63] has investigated the application of graph neural networks (GNNs) to address these challenges. In this framework, the objective is to jointly optimize the antenna placement and power allocation to maximize the system’s energy efficiency, subject to transmit power constraints and minimum inter-antenna spacing. To overcome the high computational complexity faced by conventional optimization methods, the GNN-based method treats the wireless system as a bipartite graph, with user and antenna nodes connected by edges that encode their spatial relationships. A bipartite GNN is then employed to learn effective mappings from the system state—comprising user and antenna positions as well as current channel conditions—to resource allocation decisions, specifically the antenna positions and the power allocated to each antenna. The learning process is supervised by training data generated from solutions obtained through conventional optimization techniques, enabling the GNN to approximate complex decision boundaries without requiring real-time, iterative optimization. Once the model has been trained, the GNN-based method can produce high-quality resource allocation policies with significantly lower computational overhead, making it attractive for practical systems with stringent latency requirements or large-scale user deployments.
@@
Problem (22) is challenging to solve due to the intricate coupling between the power allocation coefficients and the effective channel gains, which are nonlinear functions of the pinching-antenna positions. To address this, a practical solution is to adopt a BCD approach, which decomposes the original problem into more tractable subproblems that can be solved iteratively [54]. Specifically, the optimization proceeds as follows:
  • Power allocation optimization: Given the pinching-antenna positions, the optimal power allocation coefficients $(\alpha_{p},\alpha_{s})$ can be derived in closed-form based on the analysis in [54]. Notably, if the optimal pinching-antenna positions are available, e.g., by a linear search, this closed-form solution provides the globally optimal power allocation. This result offers a valuable theoretical benchmark for evaluating practical suboptimal algorithms with low computational complexity. However, exhaustive search quickly becomes infeasible as the number of antennas increases, motivating the need for efficient, scalable algorithms.
@@
Figure 8: Achievable data rate of the secondary user versus different numbers of pinching-antennas with $\gamma_{p}=0.1$.
Based on the BCD-SCA algorithm described above, we examine how the number of pinching antennas impacts the achievable data rate in the NOMA transmission scheme. The results are shown in in Fig. 8, where the parameters in Table VI are used and the two users are randomly deployed in a square region with side length $D=5$ m. As seen, the data rate of the secondary user increases monotonically as the number of pinching antennas grows. This improvement stems from the increased channel diversity and enhanced effective channel gain brought by each additional antenna, which collectively lead to stronger received signals and more effective interference management for the secondary user. As a result, deploying more pinching antennas along the waveguide directly contributes to improved spectral efficiency and higher achievable data rates. These findings underscore the practical importance of properly selecting and optimizing the number of pinching antennas in system design, as it enables significant performance gains while balancing hardware and implementation costs.
@@
Spatial Diversity Gains: Deploying multiple pinching antennas enables significant spatial diversity and potential array gains, which can be exploited to improve system performance. Constructive Signal Reception: Unlike conventional MIMO systems, beamforming is not feasible as only one signal can be injected into the waveguide at a time. Instead, optimizing the pinching-antenna positions to ensure that the received signals from all antennas are constructively combined at the user is crucial for performance improvement. System Design Challenge: In the multi-antenna scenario, the system design problem becomes considerably more complex, as it requires the joint optimization of pinching-antenna positions and resource allocation. Developing scalable, low-complexity algorithms for these highly coupled, non-convex problems, especially for large-scale or real-time systems, remains an important and open research direction.
@@ III-B Multiple-Waveguide Systems
To further enhance system flexibility and spatial diversity, multiple waveguides can be deployed in pinching-antenna systems. In this subsection, we systematically discuss design approaches for such multi-waveguide configurations by focusing on two representative scenarios. The first scenario considers that each waveguide is equipped with a single pinching antenna, while the second one explores the use of multiple pinching antennas per waveguide. In the remainder of this subsection, we present the signal models, formulate the corresponding optimization problems, and highlight representative solutions for each case.
@@ III-B1 Single Pinching Antenna Per Waveguide
where $e_{m}$ denotes the mean square error of estimating the information of user $m$, while ${\bf w}=[w_{1},\ldots,w_{M}]^{\top}$ are the minimum mean square error (MMSE) weight. Compared to problem (24), the WMMSE reformulation offers a more tractable structure: the objective function is convex with respect to the beamforming vectors ${\bf v}_{m}$ and the MMSE weights, though not necessarily in the antenna positions $\tilde{x}_{n}$. In contrast, the original problem contains fractional forms involving the channel vectors and beamformers, leading to a highly coupled and challenging optimization problem.
Because the optimization variables are decoupled in the constraints, the WMMSE problem can be iteratively solved by the BCD approach. Nevertheless, as highlighted in [41], optimizing the pinching-antenna positions remains non-convex, and a linear search is typically needed, which can result in substantial computational overhead for systems with long waveguides or high-frequency operation. To mitigate this issue, a low-complexity two-stage algorithm was proposed in [41]. In the first stage, an MRT-based method is used to efficiently determine the pinching antenna positions. With these locations fixed, the second stage applies the standard WMMSE algorithm to optimize the beamforming vector. This two-stage strategy circumvents the need to perform a linear search over the antenna positions in each iteration of the WMMSE algorithm, thereby greatly reducing the overall computational complexity while maintaining near-optimal performance.
@@
This setting is more challenging than the standard sum-rate maximization problem, due to the introduction of additional QoS requirements and more complex coupling between variables. Traditional BCD-type methods can become computationally prohibitive as network size and constraint set grow. To tackle this challenge, a gradient-based meta-learning joint optimization algorithm was proposed in [62]. The core idea is to “learn to optimize”: two neural networks are trained, one to update the beamforming vectors ${\bf v}_{m}$} and another one to update the pinching-antenna positions $\tilde{{\bf x}}$, by mimicking a gradient-based optimization process across a diverse set of channel realizations and QoS scenarios. During training, the networks are optimized to minimize the average negative weighted sum-rate, with the power and minimum rate constraints incorporated via penalty terms in the loss function. Once trained, the meta-learner can rapidly produce near-optimal beamforming and antenna placement solutions for new, unseen system configurations, yielding significant improvements in computational efficiency and scalability. Numerical results demonstrate that the meta-learning approach achieves better weighted sum-rate performance and rapid convergence compared to classical methods, particularly in scenarios with stringent QoS requirements. These findings indicate that data-driven and meta-learning-based algorithms are promising to tackle complex, large-scale joint optimization problems for pinching-antenna systems.
@@ III-B2 Multiple Pinching Antennas Per Waveguide
$h_{p,m}=\sum_{n=1}^{N}\frac{\eta^{\tfrac{1}{2}}e^{-j\big(\frac{2\pi}{\lambda}[(\tilde{x}_{p,n}-\bar{x})^{2}+C_{m,p}]+\frac{2\pi}{\lambda_{g}}\tilde{x}_{p,n}\big)}}{\big[(\tilde{x}_{p,n}-\bar{x})^{2}+C_{m,p}\big]^{-\frac{1}{2}}},$ (27)
@@
Based on the channel model in (27), several recent studies have investigated pinching-antenna-enabled MIMO communications [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. Specifically, the authors of [73] investigated energy-efficient pinching-antenna system designs by jointly optimizing the pinching-antenna positions and downlink beamformers. To solve the challenging optimization problem, an iterative algorithm was proposed based the BCD and SCA methods. The study demonstrates substantial transmit power savings compared to conventional massive MIMO systems. A staged GNN architecture has been proposed in [75], which efficiently learns both antenna placement and beamforming, enabling scalable and accurate resource allocation for pinching-antenna systems with strong generalization across diverse network sizes and user distributions. Meanwhile, the authors of [76] introduced a Karush-Kuhn-Tucker (KKT)-guided dual learning framework based on a transformer architecture, enabling rapid and high-quality joint optimization of transmit and pinching beamforming in multi-user pinching-antenna systems. The authors of [85] explored the use of pinching-antenna systems for uplink and downlink multi-user MIMO communication. The digital precoder and pinching-antenna positions are jointly optimized to maximize the weighted sum-rate. Two low-complexity iterative algorithms are proposed based on fractional programming (FP) and Gauss-Seidel methods. The results show that pinching-antenna systems offer significant throughput gains over conventional massive MIMO and hybrid analog-digital systems, highlighting their promise for next-generation wireless networks. A recent study [86] investigated a practical implementation of pinching-antenna systems in multi-waveguide scenarios, where antennas are installed at pre-configured discrete locations rather than being continuously adjustable. Focusing on downlink NOMA transmission, the authors formulated a joint optimization problem involving waveguide assignment, antenna activation, SIC decoding order, and power allocation. By modeling assignment and activation as coalition-formation games and applying monotonic optimization and SCA, both globally optimal and near-optimal low-complexity solutions are obtained.
@@
Enhanced Spatial Multiplexing: Leveraging multiple waveguides provides greater degrees of freedom for pinching antenna deployment, substantially boosting spatial diversity, user capacity, and spectral efficiency compared to single-waveguide or fixed-antenna systems. System Design Challenges and Solutions: Achieving optimal performance requires joint optimization of pinching-antenna positions and beamforming, resulting in highly coupled and non-convex problems. Iterative algorithms, such as BCD and WMMSE/FP-based methods, are effective in tackling these challenges. Nevertheless, developing more practical and scalable algorithms for real-time and large-scale deployment remains an important research direction. Pinching Antenna Activation: In addition to pinching-antenna optimization, efficient antenna activation, i.e., activating pinching antennas at pre-configured positions is a practical and important research direction, which enables flexible system deployment while reducing implementation complexity.
Table VII provides a structured summary of representative system design problems and corresponding design strategies for pinching-antenna systems. It highlights the diversity of architectures, optimization variables, and transmission schemes addressed in the literature. This comparative overview underscores the adaptability of the pinching-antenna framework, the progression from closed-form designs to advanced iterative and learning-based algorithms, and the importance of system complexity and application context in shaping the choice of solution strategy. These foundations set the stage for addressing more challenging and practical issues in future research on generalized pinching-antenna systems.
TABLE VII: Summary of representative system design problems and solutions for pinching-antenna systems.